# Trace field degrees in the Torelli group

### v1

@@ Abstract.
We show that for $g\geq 2$, all integers $31\leq d\leq 3g-$ 3 arise as trace field degrees of pseudo-Anosov mapping classes in the Torelli group of the closed orientable surface of genus $g$. Our method uses the Thurston-Veech construction of pseudo-Anosov maps, and we provide examples where the stretch factor has algebraic degree any even number between two and $66g-$ 6. This validates a claim by Thurston from the 1980s.
@@ 1. Introduction
Let $S_{g}$ be the closed orientable surface of genus $g\geq 2$. A homeomorphism $f$ of $S_{g}$ is pseudo-Anosov if there exists a pair of transverse singular measured foliations $\mathcal{F}^{u}$ and $\mathcal{F}^{s}$ and a real number $\lambda>1$ such that $f(\mathcal{F}^{u})=\lambda\mathcal{F}^{u}$ and $f(\mathcal{F}^{s})=\lambda^{-1}\mathcal{F}^{s}$. The number $\lambda>1$ is the stretch factor of the pseudo-Anosov map, and is an algebraic integer. The degree of the field extension $\mathbb{Q}(\lambda):\mathbb{Q}$ is the stretch factor degree, and is bounded from above by the dimension of the Teichmüller space of $S_{g}$, namely $66g-$ 6 [Thu88].
There is another field which plays a central role in this article: $\mathbb{Q}(\lambda+\lambda^{-1})$. This field is the trace field and is uniquely determined by the pair $(\mathcal{F}^{u},\mathcal{F}^{s})$ [KS00, GJ00]. The degree of the field extension $\mathbb{Q}(\lambda+\lambda^{-1}):\mathbb{Q}$ is the trace field degree, and is bounded from above by $33g-$ 3.
Strenner showed that for $S_{g}$, all integers $31\leq d\leq 3g-$ 3 arise as trace field degrees of pseudo-Anosov maps [Str17]. Furthermore, Strenner determined the set of integers which arise as stretch factor degrees: all even integers between $2$ and $\leq 6g-6$, as well as all odd integers between $3$ and $33g-$ 3 [Str17].
@@ 1.1. Torelli groups
Let $g\geq 2$. Every integer $31\leq d\leq 3g-$ 3 arises as the trace field degrees of a pseudo-Anosov mapping class in the Torelli group $\mathcal{I}(S_{g})$.
@@
Let $g\geq 2$. Every even integer $62\leq 2d\leq 6g-$ 6 arises as the stretch factor degree of a pseudo-Anosov mapping class in the Torelli group $\mathcal{I}(S_{g})$.
@@ 1.2. The Thurston–Veech construction
Given a pair of filling multicurves $\alpha,\beta\subset S_{g}$, the Thurston–Veech construction provides pseudo-Anosov mapping classes in the subgroup $\langle T_{\alpha},T_{\beta}\rangle$ of $\mathrm{Mod}(S_{g})$ generated by multitwists along the multicurves $\alpha$ and $\beta$. In his seminal 1988 Bulletin paper [Thu88], Thurston provides the upper bound of $66g-$ 6 on the algebraic degree of a pseudo-Anosov stretch factor $\lambda(f)$ and claims, without proof, that “the examples of [Thu88, Theorem 7] show that this bound is sharp”. The referenced examples are exactly the pseudo-Anosov maps in $\langle T_{\alpha},T_{\beta}\rangle$.
@@
Let $g\geq 2$ and $31\leq d\leq 3g-$ 3 be integers. Then there exists a pseudo-Anosov map on $S_{g}$ arising from the Thurston–Veech construction with trace field degree $d$ and stretch factor degree $2d$. For $g\geq 3$, the pseudo-Anosov maps can be chosen in the Torelli group $\mathcal{I}(S_{g})$.
@@ 1.3. Proof strategy for 3
For a pair $\alpha,\beta\subset S_{g}$ of filling multicurves, let $X=\left(|\alpha_{i}\cap\beta_{j}|\right)_{ij}$ be the matrix encoding the number of intersections of the components of $\alpha$ and $\beta$.
The matrix $XX^{\top}$ is primitive, hence by Perron-Frobenius theory its spectral radius equals its largest eigenvalue and is therefore an algebraic integer. Let $d$ be its algebraic degree. We call the number $d$ the multicurve intersection degree of $\alpha$ and $\beta$.
@@
Assuming 4, what remains to be done in order to prove the first part of 3 is to construct all multicurve intersection degrees $31\leq d\leq 3g-$ 3 on $S_{g}$ for $g\geq 2$. By the Thurston–Veech construction, the trace field degree of the resulting pseudo-Anosov maps equals exactly the multicurve intersection degree of the multicurves $\alpha$ and $\beta$ used in the construction [Thu88, Vee89]. Hence, we are done by setting $\varepsilon=1$ in 4.
In order to prove the Torelli part of 3, we construct the multicurves $\alpha$ and $\beta$ realising the multicurve intersection degrees $31\leq d\leq 3g-$ 3 in such a way that $T_{\alpha}\circ T_{\beta}^{-1}$ is an element of $\mathcal{I}(S_{g})$. We will ensure this by choosing multicurves $\alpha$ and $\beta$ which consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. We can then finish the proof of 3 by setting $\varepsilon=-1$ in 4. To see this, note that if $T_{\alpha}\circ T_{\beta}^{-1}$ is an element of $\mathcal{I}(S_{g})$, then so is $T_{\alpha}^{n}\circ T_{\beta}^{-n}$.
We note that our examples of multicurves $\alpha$ and $\beta$ that we use in order to construct examples in the Torelli group $\mathcal{I}(S_{g})$ cannot yield multicurve degrees greater than one on the surface $S_{2}$. Indeed, there exist no bounding pairs on $S_{2}$ and a multicurve can have at most one separating component.
To find the suitable multicurves $\alpha$ and $\beta$ realising all possible multicurve intersection degrees $31\leq d\leq 3g-$ 3 is the main technical contribution in this article.
@@ 1.4. Proof of 1 and 2 for $g=2$
pseudo–Anosov mapping class $[f]\in\mathrm{Mod}(S_{2,1})$ minimal polynomial of $\lambda(f)$
@@
In order to obtain elements in $\mathrm{Mod}(S_{2})$, we appeal to the forgetful map. Since our examples do not have $1$-pronged singularity at the puncture, we can fill in it in order to get a pseudo–Anosov mapping class, with the same stretch factor, in $\mathrm{Mod}(S_{2})$, see [HK06, Lemma 2.6] for details. This completes the proof for $g=2$.
@@ 1.5. Explicit examples
More precisely, we provide an almost explicit construction of multicurves with intersection degrees $31<d\leq 3g-$ 3 for which the pseudo-Anosov mapping class $T_{\alpha}\circ T_{\beta}$ has stretch factor degree $2d$. In the setting of 4, this means for $\varepsilon=1$ one can choose $n=1$. Our argument uses [LL24, Theorem 6] (see Section 6 for details). Unfortunately, for the multicurves we construct in the Torelli case, the criterion from [LL24] does not apply. Hence, our need to use 4 for the proof of 3, which provides a slightly less explicit result.
Aided by the computer, we can subsequently find completely explicit examples of multicurves and therefore entirely explicit pseudo-Anosov mapping classes on $S_{g}$ arising from the Thurston–Veech construction realising the maximal stretch factor degree $66g-$ 6, for all genera $g\leq 201$.
@@
While 4 provides the existence of field extensions $\mathbb{Q}(\lambda):\mathbb{Q}(\lambda+\lambda^{-1})$ of degree two for mapping classes in $\langle T_{\alpha},T_{\beta}\rangle$, realising extensions of degree one seems to be more mysterious. For example, Veech [Vee82] discovered a family of Hecke groups $\langle T_{\alpha},T_{\beta}\rangle=\langle\left(\begin{smallmatrix}1&\lambda_{q}\\ 0&1\end{smallmatrix}\right),\left(\begin{smallmatrix}1&0\\ -\lambda_{q}&1\end{smallmatrix}\right)\rangle$, where $\lambda_{q}=2\cos\pi/q$ for $q\geq 3$. The genus of the surface $S_{g}$ is $(q-1)/2$ for odd $q$. For $q=7,9$ one can find stretch factors of degree one over the trace field $\mathbb{Q}(\lambda_{q})$: for instance $T_{\alpha}\circ T_{\beta}^{-1}$ is an example for $q=7$, and we refer to [Bou22] for $q=9$. However, it is conjectured (see [HMTY08, Remark 9]) that stretch factors of degree one over $\mathbb{Q}(\lambda_{q})$ do not exist for odd $q\geq 11$.
@@ 1.8. A natural field extension from the perspective of curves
A pair of filling multicurves $\alpha,\beta\subset S_{g}$ naturally determines a bipartite graph whose vertices correspond to curve components and the number of edges between each pair of vertices equals the number of intersection points of the respective curve components. The adjacency matrix of this graph is $\Omega=\left(\begin{smallmatrix}0&X\\ X^{\top}&0\end{smallmatrix}\right)$. Clearly, the square root $\sqrt{\mu}$ of the spectral radius $\mu$ of $XX^{\top}$ equals the spectral radius of $\Omega$. We call the algebraic degree of $\sqrt{\mu}$ the multicurve bipartite degree of $\alpha$ and $\beta$. Similarly to the field extension $\mathbb{Q}(\lambda):\mathbb{Q}(\lambda+\lambda^{-1})$, also the field extension $\mathbb{Q}(\sqrt{\mu}):\mathbb{Q}(\mu)=\mathbb{Q}(\lambda+\lambda^{-1})$ has degree one or two. We prove the following result.
@@
Every even integer $62\leq 2d\leq 6g-$ 6 is realised as a multicurve bipartite degree on $S_{g}$ for $g\geq 2$.
@@ Acknowledgments
The authors thank Dan Margalit and Jean-Claude Picaud for inspiring discussions, Mark Bell for his help with the software flipper [Bel21] and Curt McMullen for comments on an earlier version of the article. They are particularly grateful to Dan Margalit for pointing out a mistake in the formulation of the proof of 3 in an earlier version of the article. The first author has been partially supported by the LabEx PERSYVAL-Lab (ANR-11-LABX-0025-01) funded by the French program Investissement d’avenir. The first author would like to thank the University of Fribourg and the Center for Mathematical Modeling (CMM) at Universidad de Chile for excellent working conditions.
@@ 2. A nonsplitting criterion
Since $\lambda=\lambda(f)$ solves the quadratic equation $t^{2}-(\lambda+\lambda^{-1})t+1=0$$\lambda$ has degree $1$ or $2$ over $K$. All what we need to do is find $n\in\mathbb{Z}_{>0}$ such that $\lambda\not\in K$, or equivalently such that the discriminant $D=(2-\varepsilon(2nr)^{2})^{2}-4=16\cdot n^{2}\cdot((n\varepsilon\mu)^{2}-\varepsilon\mu)$ of the quadratic equation is not a square in $K$. We will proceed by contradiction. Let $\mu^{\prime}=\varepsilon\mu$ and let us assume that $(n\mu^{\prime})^{2}-\mu^{\prime}$ is a square in $K=\mathbb{Q}(\mu^{\prime})$ for every $n>0$. Since the expression is invariant under the transformation $n\mapsto-n$, we can assume the expression is a square for every $n\in\mathbb{Z}\setminus\{0\}$.
Let $P=a_{d}t^{d}+a_{d-1}t^{d-1}+\dots+a_{1}t+a_{0}\in\mathbb{Z}[t]$ be the minimal polynomial of $\mu^{\prime}$ over $\mathbb{Q}$. The Thurston–Veech construction implies that $\mu$ is an eigenvalue of a square matrix, so is $\mu^{\prime}$ and $a_{d}=1$. Thus, $\mu^{\prime}$ and $n^{2}\mu^{\prime}-1$ are algebraic units. The norm of $\mu^{\prime}$ equals $N(\mu^{\prime})=(-1)^{d}a_{0}$. Similarly, the minimal polynomial of $n^{2}\mu^{\prime}-1$ is $n^{2d}P\left(\frac{t+1}{n^{2}}\right)$. Inspecting the constant term, we have
@@
Altogether this gives $N((n\mu^{\prime})^{2}-\mu^{\prime})=Q(n^{2})$, where
@@
By assumption, $Q(n^{2})$ is a square for every $n\in\mathbb{Z}\setminus\{0\}$. We show that $Q(0)=N(-\mu^{\prime})$ is also a square. Indeed, for any prime integer $p$, the reduction modulo $n=p$ of $N((n\mu^{\prime})^{2}-\mu^{\prime})$ gives that $N(-\mu^{\prime})$ is a quadratic residue. Thus it is also a square in $\mathbb{Z}$. Hence $Q(t)\in\mathbb{Z}[t]$ is a polynomial taking integral square value at every integer specialisation. By a result of Murty [Mur08, Theorem 1]$Q(t^{2})$ is the square of a polynomial.
Moreover, we observe that $Q(t)=a_{0}\cdot t^{d}P\left(\frac{1}{t}\right)\in\mathbb{Q}[t]$. In particular $Q\left(\frac{1}{\mu^{\prime}}\right)=0$, and since $\mu^{\prime}$ and $\frac{1}{\mu^{\prime}}$ generate the field extension $K:\mathbb{Q}$, the polynomial $Q$ is irreducible over $\mathbb{Q}$. It is in particular separable. Now each root $0\neq a\in\mathbb{C}$ of $Q$ gives rise to two distinct roots ±$\pm\sqrt{a}$ of $Q(t^{2})$, and conversely. Thus $Q(t^{2})$ is also separable, and cannot be a square. This concludes the proof of the theorem. ∎

### v2

@@ Abstract.
We show that for $g\geq 2$, all integers $1\leq d\leq 3g-3$ arise as trace field degrees of pseudo-Anosov mapping classes in the Torelli group of the closed orientable surface of genus $g$. Our method uses the Thurston–Veech construction of pseudo-Anosov maps, and we provide examples where the stretch factor has algebraic degree any even number between two and 6g66g-6. This validates a claim by Thurston from the 1980s.
@@ 1. Introduction
Let $S_{g}$ be the closed orientable surface of genus $g\geq 2$. A homeomorphism $f$ of $S_{g}$ is pseudo-Anosov if there exists a pair of transverse singular measured foliations $\mathcal{F}^{u}$ and $\mathcal{F}^{s}$ and a real number $\lambda>1$ such that $f(\mathcal{F}^{u})=\lambda\mathcal{F}^{u}$ and $f(\mathcal{F}^{s})=\lambda^{-1}\mathcal{F}^{s}$. The number $\lambda>1$ is the stretch factor of the pseudo-Anosov map, and is an algebraic integer. The degree of the field extension $\mathbb{Q}(\lambda):\mathbb{Q}$ is the stretch factor degree, and is bounded from above by the dimension of the Teichmüller space of $S_{g}$, namely 6g66g-6 [Thu88].
There is another field which plays a central role in this article: $\mathbb{Q}(\lambda+\lambda^{-1})$. This field is the trace field and is uniquely determined by the pair $(\mathcal{F}^{u},\mathcal{F}^{s})$ [KS00, GJ00]. The degree of the field extension $\mathbb{Q}(\lambda+\lambda^{-1}):\mathbb{Q}$ is the trace field degree, and is bounded from above by 3g33g-3.
Strenner showed that for $S_{g}$, all integers $1\leq d\leq 3g-3$ arise as trace field degrees of pseudo-Anosov maps [Str17]. Furthermore, Strenner determined the set of integers which arise as stretch factor degrees: all even integers between $2$ and $\leq 6g-6$, as well as all odd integers between $3$ and 3g33g-3 [Str17], but none of Strenner’s examples are elements of the Torelli group (see Section 1.3).
@@ 1.1. Torelli groups
Let $g\geq 2$. Every integer $1\leq d\leq 3g-3$ arises as the trace field degrees of a pseudo-Anosov mapping class in the Torelli group $\mathcal{I}(S_{g})$.
@@
Let $g\geq 2$. Every even integer $2d$ with $2\leq 2d\leq 6g-6$ arises as the stretch factor degree of a pseudo-Anosov mapping class in the Torelli group $\mathcal{I}(S_{g})$.
@@ 1.2. The Thurston–Veech construction
Given a pair of filling multicurves $\alpha,\beta\subset S_{g}$, the Thurston–Veech construction provides pseudo-Anosov mapping classes in the subgroup $\langle T_{\alpha},T_{\beta}\rangle$ of $\mathrm{Mod}(S_{g})$ generated by multitwists along the multicurves $\alpha$ and $\beta$. In his seminal 1988 Bulletin paper [Thu88], Thurston provides the upper bound of 6g66g-6 on the algebraic degree of a pseudo-Anosov stretch factor $\lambda(f)$ and claims, without proof, that “the examples of [Thu88, Theorem 7] show that this bound is sharp”. The referenced examples are exactly the pseudo-Anosov maps in $\langle T_{\alpha},T_{\beta}\rangle$.
@@
Let $g\geq 2$ and $1\leq d\leq 3g-3$ be integers. Then there exists a pseudo-Anosov map on $S_{g}$ arising from the Thurston–Veech construction with trace field degree $d$ and stretch factor degree $2d$. For $g\geq 3$, the pseudo-Anosov maps can be chosen in the Torelli group $\mathcal{I}(S_{g})$.
@@ 1.3. Proof strategy for 3
For a pair $\alpha,\beta\subset S_{g}$ of filling multicurves, let $X=\left(|\alpha_{i}\cap\beta_{j}|\right)_{ij}$ be the matrix encoding the number of geometric intersections of the components of $\alpha$ and $\beta$.
The matrix $XX^{\top}$ is primitive, hence by Perron–Frobenius theory its spectral radius equals its largest eigenvalue and is therefore an algebraic integer. Let $d$ be its algebraic degree. We call the number $d$ the multicurve intersection degree of $\alpha$ and $\beta$.
@@
Assuming 4, what remains to be done in order to prove the first part of 3 is to construct all multicurve intersection degrees $1\leq d\leq 3g-3$ on $S_{g}$ for $g\geq 2$. By the Thurston–Veech construction, the trace field degree of the resulting pseudo-Anosov maps equals exactly the multicurve intersection degree of the multicurves $\alpha$ and $\beta$ used in the construction [Thu88, Vee89]. Hence, we are done by setting $\varepsilon=1$ in 4.
In order to prove the Torelli part of 3, we construct the multicurves $\alpha$ and $\beta$ realising the multicurve intersection degrees $1\leq d\leq 3g-3$ in such a way that $T_{\alpha}\circ T_{\beta}^{-1}$ is an element of $\mathcal{I}(S_{g})$. We will ensure this by choosing multicurves $\alpha$ and $\beta$ which consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. We can then finish the proof of 3 by setting $\varepsilon=-1$ in 4. To see this, note that if $T_{\alpha}\circ T_{\beta}^{-1}$ is an element of $\mathcal{I}(S_{g})$, then so is $T_{\alpha}^{n}\circ T_{\beta}^{-n}$.
We note that our examples of multicurves $\alpha$ and $\beta$ that we use in order to construct examples in the Torelli group $\mathcal{I}(S_{g})$ cannot yield multicurve intersection degrees greater than one on the surface $S_{2}$. Indeed, there exist no bounding pairs on $S_{2}$ and a multicurve can have at most one separating component.
The constructed examples in [Str17] do not belong to the Torelli group. Indeed, the first components of the considered multicurves $A$ and $B$ [Str17, Figure 6.1] are nonzero in homology and they intersect nontrivially.
To find the suitable multicurves $\alpha$ and $\beta$ realising all possible multicurve intersection degrees $1\leq d\leq 3g-3$ is the main technical contribution in this article.
@@ 1.4. Proof of 1 and 2 for $g=2$
pseudo–Anosov mapping class $f]\in\mathrm{Mod}(S_{2,1})$ minimal polynomial of $\lambda(f)$
@@
In order to obtain elements in $\mathrm{Mod}(S_{2})$, we appeal to the forgetful map. Since our examples do not have a $1$-pronged singularity at the puncture, we can fill it in order to get a pseudo-Anosov mapping class, with the same stretch factor, in $\mathrm{Mod}(S_{2})$, see [HK06, Lemma 2.6] for details. This completes the proof of 2 for $g=2$. Now for each example, the minimal polynomial of $\lambda(f)$ is reciprocal and has degree $2d$ (for $d=1,2,3$ respectively). Hence the minimal polynomial of $\lambda(f)+\lambda(f)^{-1}$ has degree $d$ as required, and 1 is proved.
@@ 1.5. Explicit examples
More precisely, we first provide an almost explicit construction of multicurves with intersection degrees $1<d\leq 3g-3$ for which the pseudo-Anosov mapping class $T_{\alpha}\circ T_{\beta}$ has stretch factor degree $2d$. In the setting of 4, this means for $\varepsilon=1$ one can choose $n=1$. Our argument uses [LL24, Theorem 6] (see Section 6 for details). Unfortunately, for the multicurves we construct in the Torelli case, the criterion from [LL24] does not apply. Hence our need to use 4 for the proof of 3, which provides a slightly less explicit result.
Aided by the computer, we can subsequently find completely explicit examples of multicurves and therefore entirely explicit pseudo-Anosov mapping classes on $S_{g}$ arising from the Thurston–Veech construction realising the maximal stretch factor degree 6g66g-6, for all genera $g\leq 201$.
We emphasise that this construction, contained in Section 6, is independent of the proof of our main results, 1, 2 and 3.
@@
While 4 provides the existence of field extensions $\mathbb{Q}(\lambda):\mathbb{Q}(\lambda+\lambda^{-1})$ of degree two for mapping classes in $\langle T_{\alpha},T_{\beta}\rangle$, realising extensions of degree one seems to be more mysterious. For example, Veech [Vee82] discovered a family of Hecke groups $\langle T_{\alpha},T_{\beta}\rangle=\langle\left(\begin{smallmatrix}1&\lambda_{q}\\ 0&1\end{smallmatrix}\right),\left(\begin{smallmatrix}1&0\\ -\lambda_{q}&1\end{smallmatrix}\right)\rangle$, where $\lambda_{q}=2\cos\pi/q$ for $q\geq 3$. The genus of the surface $S_{g}$ is (q1)/2(q-1)/2 for odd $q$. For $q=7,9$ one can find stretch factors of degree one over the trace field $\mathbb{Q}(\lambda_{q})$: for instance $T_{\alpha}\circ T_{\beta}^{-1}$ is an example for $q=7$, and we refer to [Bou22] for $q=9$. However, it is conjectured (see [HMTY08, Remark 9]) that stretch factors of degree one over $\mathbb{Q}(\lambda_{q})$ do not exist for odd $q\geq 11$.
@@ 1.8. A natural field extension from the perspective of curves
A pair of filling multicurves $\alpha,\beta\subset S_{g}$ naturally determines a bipartite graph whose vertices correspond to curve components and the number of edges between each pair of vertices equals the number of intersection points of the respective curve components. The adjacency matrix of this graph is $\Omega=\left(\begin{smallmatrix}0&X\\ X^{\top}&0\end{smallmatrix}\right)$. Clearly, the square root $\sqrt{\mu}$ of the spectral radius $\mu$ of $XX^{\top}$ equals the spectral radius of $\Omega$. We call the algebraic degree of $\sqrt{\mu}$ the multicurve bipartite degree of $\alpha$ and $\beta$. Similarly to the field extension $\mathbb{Q}(\lambda):\mathbb{Q}(\lambda+\lambda^{-1})$, also the field extension $\mathbb{Q}(\sqrt{\mu}):\mathbb{Q}(\mu)$ has degree one or two. From the point of view of the Thurston–Veech construction one has $\lambda(f)+\lambda(f)^{-1}=2+\mu$ for $f=T_{\alpha}\circ T_{\beta}^{-1}$ (see the proof of 4). We prove the following result.
@@
Every even integer $2\leq 2d\leq 6g-6$ is realised as a multicurve bipartite degree on $S_{g}$ for $g\geq 2$.
@@ Acknowledgments
The authors thank Dan Margalit and Jean-Claude Picaud for inspiring discussions, Mark Bell for his help with the software flipper [Bel21] and Curt McMullen for comments on an earlier version of the article. They are particularly grateful to Dan Margalit for pointing out a mistake in the formulation of the proof of 3 in an earlier version of the article. The authors also thank the anonymous referee for several useful suggestions, comments and corrections. The first author has been partially supported by the LabEx PERSYVAL-Lab (ANR-11-LABX-0025-01) funded by the French program Investissement d’avenir. The first author would like to thank the University of Fribourg and the Center for Mathematical Modeling (CMM) at Universidad de Chile for excellent working conditions.
@@ 2. A nonsplitting criterion
We need the following lemma which is a small extension of a result of Murty [Mur08, Theorem 4]. It is probably well-known but we could not find a reference so we include a proof here.
Lemma 6.
Let $Q(t)\in\mathbb{Z}[t]$ be a polynomial taking an integral square value at all but finitely many positive integer specialisations. Then $Q(t)$ is the square of a polynomial in $\mathbb{Z}[t]$.
Proof.
Let $N\in\mathbb{N}$ be a positive integer such that $Q(n)$ is a square for every $n\in\mathbb{N}$ with $n>N$. Let $n\in\mathbb{Z}$. Consider a prime number $p$ such that $n+p>N$. The integer $Q(n+p)$ is a square by assumption, and its reduction modulo $p$ is $Q(n)$. Thus, $Q(n)\in\mathbb{Z}$ is a quadratic residue for all but finitely many prime numbers. So the value $Q(n)$ is a square in $\mathbb{Z}$ (see [Hal33, Theorem 1 (end of the proof)]). Thus we can apply [Mur08, Theorem 4] to conclude that $Q(t)$ is the square of a polynomial in $\mathbb{Z}[t]$. ∎
@@
Since $\lambda=\lambda(f)$ solves the quadratic equation $t^{2}-(\lambda+\lambda^{-1})t+1=0$$\lambda$ has degree $1$ or $2$ over $K$. All that we need to do is find $n\in\mathbb{Z}_{>0}$ such that $\lambda\not\in K$, or equivalently such that the discriminant $D=(2-\varepsilon(2nr)^{2})^{2}-4=16\cdot n^{2}\cdot((n\varepsilon\mu)^{2}-\varepsilon\mu)$ of the quadratic equation is not a square in $K$. We proceed by contradiction. Let $\mu^{\prime}=\varepsilon\mu$ and let us assume that $(n\mu^{\prime})^{2}-\mu^{\prime}$ is a square in $K=\mathbb{Q}(\mu^{\prime})$ for every $n>0$. Since the expression $(n\mu^{\prime})^{2}-\mu^{\prime}$ is invariant under the transformation $n\mapsto-n$, we can assume that $(n\mu^{\prime})^{2}-\mu^{\prime}$ is a square for every $n\in\mathbb{Z}\setminus\{0\}$.
Let $q\in\mathbb{Q}[t]$ be the minimal polynomial of $\frac{1}{\mu^{\prime}}$ over $\mathbb{Q}$. Our strategy is to use 6 to show the following claim.
Claim. The polynomial $q(t^{2})\in\mathbb{Q}[t]$ has a double root in $\mathbb{C}$.
This claim directly leads to a contradiction. Indeed, the polynomial $q\in\mathbb{Q}[t]$ is irreducible, and thus only has simple roots. Now each root $0\neq a\in\mathbb{C}$ of $q$ gives rise to two distinct roots $\pm\sqrt{a}$ of $q(t^{2})$, contradicting the claim. In order to finish the proof of 4, it therefore remains to prove the claim.
Proof of Claim. Let $P=a_{d}t^{d}+a_{d-1}t^{d-1}+\dots+a_{1}t+a_{0}\in\mathbb{Q}[t]$ be the minimal polynomial of $\mu^{\prime}$ over $\mathbb{Q}$ with $a_{d}=1$. The Thurston–Veech construction implies that $\mu$ is an eigenvalue of a square matrix with integer coefficients, so $\mu^{\prime}$ is also an eigenvalue of a square matrix with integer coefficients, implying that $P\in\mathbb{Z}[t]$.
Thus, $\mu^{\prime}$ and $n^{2}\mu^{\prime}-1$ are algebraic integers, and we can consider their norms $N(\cdot)$. Obviously, for $\mu^{\prime}$ one has $N(\mu^{\prime})=(-1)^{d}a_{0}$. Let us compute the norm of $n^{2}\mu^{\prime}-1$. The monic polynomial $n^{2d}P\left(\frac{t+1}{n^{2}}\right)\in\mathbb{Z}[t]$ has degree $d$ and admits $n^{2}\mu^{\prime}-1$ as a root. Moreover the fields $\mathbb{Q}(\mu^{\prime})$ and $\mathbb{Q}(n^{2}\mu^{\prime}-1)$ coincide, so they have the same degree over $\mathbb{Q}$, namely $d$. Hence $n^{2d}P\left(\frac{t+1}{n^{2}}\right)$ is the minimal polynomial of $n^{2}\mu^{\prime}-1$. Inspecting its constant term, we deduce
@@
Since the norm is multiplicative, we have
$N((n\mu^{\prime})^{2}-\mu^{\prime})=N(n^{2}\mu^{\prime}-1)\cdot N(\mu^{\prime})=a_{0}\sum_{k=0}^{d}a_{k}n^{2d-2k}=Q(n^{2}),$
where
@@
Since by assumption $(n\mu^{\prime})^{2}-\mu^{\prime}$ is a square for every $n\in\mathbb{Z}\setminus\{0\}$, we deduce that $Q(n^{2})$ is a square for every $n\in\mathbb{Z}\setminus\{0\}$. By 6, $Q(t^{2})$ is itself a square in $\mathbb{Z}[t]$.
To finish the proof, observe that $Q(t)=a_{0}\cdot t^{d}P\left(\frac{1}{t}\right)\in{\mathbb{Z}}[t]$. In particular, $Q\left(\frac{1}{\mu^{\prime}}\right)=0$ and $Q$ has degree $d$ over $\mathbb{Q}$. Since $\mu^{\prime}$ and $\frac{1}{\mu^{\prime}}$ generate the same field extension $K:\mathbb{Q}$, the minimal polynomial $q$ of $\frac{1}{\mu^{\prime}}$ over $\mathbb{Q}$ is also of degree $d$. This implies that $Q$ must be a rational multiple of $q$. In particular, $q(t^{2})$ is a rational multiple of $Q(t^{2})$ and therefore has a double root in $\mathbb{C}$. ∎